OpenStudy (anonymous):

The expression (a^-1) - (b^-1)/ (a^-2)- (b^-2) (a) b-a/ ab (b) b+a/ ab (c) ab/ b+a (d) (a^2)-(b^2)/ a-b (e) a-b

OpenStudy (callisto):

(a^-1) - (b^-1)/ (a^-2)- (b^-2) \[=\frac{\frac{1}{a} - \frac{1}{b}}{\frac{1}{a^2}- \frac{1}{b^2}}\]\[=\frac{\frac{b}{ab} - \frac{a}{ab}}{\frac{b^2}{a^2b^2}- \frac{a^2}{a^2b^2}}\]\[=\frac{\frac{b-a}{ab}}{\frac{b^2-a^2}{a^2b^2}}\]\[=\frac{b-a}{ab} \times \frac{a^2b^2}{b^2-a^2}\]\[=\frac{b-a}{ab} \times \frac{a^2b^2}{(b-a)(b+a)}\]\[=...?\]

OpenStudy (anonymous):

i got (d) is that right?

OpenStudy (callisto):

Not really.... Do you understand the above workings?

OpenStudy (anonymous):

yeah. you're supposed to multiply the numerators and the denominators. i got (a^2b^3)-(a^3b^2)/ (ab^3)-(a^3b).

OpenStudy (anonymous):

and then i simplified it down to a^2 - b^2/ a-b

OpenStudy (callisto):

So weird... Actually, all you need to do is to cancel the common factors there :|

OpenStudy (anonymous):

oh and then you get (a) right?

OpenStudy (anonymous):

wait nevermind

OpenStudy (anonymous):

i'm confused is it (e)?

OpenStudy (callisto):

Not really :|

OpenStudy (anonymous):

could u help me out?

OpenStudy (callisto):

Actually, which part you don't understand?

OpenStudy (anonymous):

the cancelling of the common factors

OpenStudy (callisto):

Just a minute :|

OpenStudy (anonymous):

ok

OpenStudy (callisto):

Assume you have the expression \[\frac{xy}{z} \times \frac{xz}{y}\]